OpenStudy (anonymous):

true or false: if f is one-to-one, then f^-1(x)=1/f(x)

OpenStudy (agent0smith):

y = 1/x is a one-to-one function. It's inverse is y = 1/x. Is 1/x equal to 1/(1/x)?

OpenStudy (anonymous):

\[f ^{-1}(x)=1divf(x)\] is the equation so where is y coming from?

OpenStudy (agent0smith):

y is easier to write than f(x) when using inverses.

OpenStudy (anonymous):

you said y=1/x and it's inverse is also y=1/x

OpenStudy (agent0smith):

Correct.